The moon has a mass of 7.34 × 1022 kg and a radius of 1.74 × 106 meters. If


you have a mass of 66 kg, how strong is the force between you and the moon?

1
Expert's answer
2021-11-09T16:35:41-0500

"F_m = \frac{GmM_m}{R_m^2}"

G = gravitational constant

m = mass of the person

Mm = mass of the Moon

Rm = radius of the Moon

"F_m = \frac{6.673 \times 10^{-11} \times 66 \times 7.34 \times 10^{22}}{(1.74 \times 10^6)^2} \\

= 106.77 \;N"